OpenStudy (anonymous):

Solve for the variables. 3c - 5d = -8 -2c + 10d = 8 Question 1 options: a) c= 1/5 d= 2 b) c= -2, d= 2/5 c) c = -2, d = 1 d) None of these

OpenStudy (jonnyvonny):

You can do this one of two ways; substitution or elimination, which do you want to do?

OpenStudy (anonymous):

When I do substitution it said it was wrong.

OpenStudy (jonnyvonny):

The answer shouldn't change, so I'm assuming you did it incorrectly. We shall use the substitution method we may know what you did wrong (unless we're both wrong then cry.jpg). So, lets get the c alone in the second equation, what do you get? post answer and I shall as well.

OpenStudy (jonnyvonny):

c=5d-4

OpenStudy (anonymous):

when i substituted i got -13=-8 which dont equal

OpenStudy (jonnyvonny):

You most likely distributed the 3 incorrectly when you plugged in c for "(3c-5d)=-8"

OpenStudy (jonnyvonny):

3(5d-4)-5d=-8 15d-12-5d=-8 10d-12=-8 10d=4 d=2/5 I did the other one wrong xD

OpenStudy (jonnyvonny):

From then, we can plug in d=2/5 into any equation: 3c-5d=-8 3c-5(2/5)=-8 3c-2=-8 3c=-6 c=-3

OpenStudy (jonnyvonny):

Do you understand the process?

OpenStudy (anonymous):

Yes mam. Thank you.